# Img-Diff: Contrastive Data Synthesis for Multimodal Large Language Models

### v2

@@
Abstract.
High-performance Multimodal Large Language Models (MLLMs) rely heavily on data quality. This study introduces a novel dataset named Img-Diff, designed to enhance fine-grained image recognition in MLLMs by leveraging insights from contrastive learning and image difference captioning. By analyzing object differences between similar images, we challenge models to identify both matching and distinct components. We utilize the Stable-Diffusion-XL model and advanced image editing techniques to create pairs of similar images that highlight object replacements. Our methodology includes a Difference Area Generator for object differences identifying, followed by a Difference Captions Generator for detailed difference descriptions. The result is a relatively small but high-quality dataset of “object replacement” samples. We use the the proposed dataset to finetune state-of-the-art (SOTA) MLLMs such as MGM-7B, yielding comprehensive improvements of performance scores over SOTA models that trained with larger-scale datasets, in numerous image difference and Visual Question Answering tasks. For instance, our trained models notably surpass the SOTA models GPT-4V and Gemini on the MMVP benchmark. Besides, we investigate alternative methods for generating image difference data through “object removal” and conduct a thorough evaluation to confirm the dataset’s diversity, quality, and robustness, presenting several insights on the synthesis of such a contrastive dataset. To encourage further research and advance the field of multimodal data synthesis and enhancement of MLLMs’ fundamental capabilities for image understanding, we release our codes and dataset at https://github.com/modelscope/data-juicer/tree/ImgDiff.
copyright: none
@@ Abstract.
Figure 1. Three “object replacement” examples within Img-Diff, highlighting fine-grined difference in both vision and language.
[…]
Examples of generated “object replacement” data within Img-Diff.
1. Introduction
@@
The emergence of large language models (LLMs) has revolutionized natural language processing (Mann et al., 2020; Touvron et al., 2023), also paved the way for the development of Multimodal Large Language Models (MLLMs) that seamlessly integrate linguistic and visual understanding. Improving the performance of MLLMs hinges on two primary avenues: evolving model architectures and enhancing dataset quality (Qin et al., 2024). The majority of state-of-the-art (SOTA) MLLMs (Liu et al., 2023c, b, 2024b; Bai et al., 2023; Li et al., 2023d) employ a two-phase approach, commencing with a pretraining phase involving extensive image-text pairs for modality alignment, followed by fine-tuning aimed at optimizing visual question answering (VQA) capabilities with specific instruction tuning datasets.
@@
The efficacy of pretraining datasets profoundly affects MLLMs’ capabilities in performing core visual tasks. Concurrently, the quality of visual instruction tuning datasets plays a crucial role in the overall performance of MLLMs in VQA tasks and diverse downstream applications. With the evolution of visual instruction tuning datasets, several recent studies have explored integration of object detection and Optical Character Recognition (OCR) datasets, such as Refcoco (Kazemzadeh et al., 2014), Visual Genome (Krishna et al., 2017), OCR-VQA (Mishra et al., 2019), and TextVQA (Singh et al., 2019), effectively enhancing MLLMs’ proficiency in tasks requiring detailed image perception.
@@
In this paper, we focus on a new direction for enhancing MLLM datasets, driven by the potential of pairs of different images to refine models’ image recognition capabilities, as demonstrated by advancements in contrastive learning and image difference captioning for conventional vision models (Zhang et al., 2023b; Black et al., 2024; Park et al., 2019; Jhamtani and Berg-Kirkpatrick, 2018). Specifically, we introduce a general-purpose yet challenging dataset named Img-Diff, which sets itself apart from existing visual instruction tuning datasets of MLLMs by generating pairs of highly similar images featuring subtle object alterations. Rather than compelling MLLMs to focus solely on a single image, our dataset challenges them to analyze paired images and articulate the differences within designated regions, meanwhile taking the high-quality textual descriptions of the difference as learning signals. By doing so, we aim to empower MLLMs with enhanced capabilities of fine-grained image recognition, via scrutinizing intricate areas within images and assess discrepancies across corresponding regions.
@@
Following the construction of our dataset, we integrate it into the original visual instruction tuning datasets of LLaVA-1.5 (Liu et al., 2023b) and MGM (Li et al., 2023d). Subsequently, we perform fine-tuning on these models and evaluate their performance on image difference benchmarks, including MMVP (Tong et al., 2024), Spot-the-Diff (Jhamtani and Berg-Kirkpatrick, 2018), and Image-Edit-Request (Tan et al., 2019), as well as widely recognized MLLM benchmarks. Our evaluation results reveal that after fine-tuning with our Img-Diff dataset, LLaVA-1.5-7B and MGM-7B achieve notable enhancements in image difference benchmarks, aligning their performance with state-of-the-art (SOTA) models. For instance, they notably surpass the SOTA models GPT-4V(Achiam et al., 2023) and Gemini(Google, 2023) on the MMVP benchmark. Moreover, these models exhibit comprehensive improvements across numerous well-recognized MLLM benchmarks, achieving an average score improvement of up to 3.06%, underscoring the useful role our dataset plays in bolstering MLLMs’ competencies in both image difference recognition and fine-grained image analysis.
@@
We further evaluate the diversity and quality of our dataset, ensuring it encompasses a broad array of object categories while showcasing rich variability. Through meticulous manual labeling, we affirm the high quality of the dataset. Additionally, we conduct ablation studies to examine the effects of various filtering intensities. We also investigate an alternative methodology for constructing image difference data focused on “object removal”, assessing its effectiveness on LLaVA-1.5-7B and MGM-7B, and present fruitful insights on the construction of contrastive data synthesis.
@@
  • We present a novel data synthesis method and an effect-proven Img-Diff dataset, comprising pairs of highly similar images, with a focus on processes such as segmentation, filtering, and detailed captioning of image difference.
@@
  • We execute visual instruction tuning on LLaVA-1.5-7B and MGM-7B using our dataset, and rigorously assess the fine-tuned models’ performance on many widely-used MLLM benchmarks, demonstrating substantial performance improvements.
@@
  • We provide a comprehensive evaluation of the diversity and quality of the proposed dataset, confirming its richness and high standard. Through ablation studies, we identify good empirical filtering thresholds for such contrastive datasets.
@@
  • To facilitate ongoing research, encouraging innovative endeavors in more MLLM datasets and image difference methodologies, we open-source our dataset and codes at https://github.com/modelscope/ data-juicer/tree/ImgDiff.
2. Background and Related Works
2.1. Multimodal Large Language Models
@@
MLLMs have exhibited remarkable advancements and continue to evolve rapidly (Achiam et al., 2023; Liu et al., 2024b, a; Google, 2023). Research highlights two key factors that primarily influence the effectiveness of MLLMs: model architecture and dataset quality (Qin et al., 2024).
@@
With respect to model architecture, notable approaches include Flamingo (Alayrac et al., 2022), IDEFICS (Hugo et al., 2023; Laurençon et al., 2024), BLIP-2 (Li et al., 2023b), and Qwen-VL (Bai et al., 2023), which leverage learnable queries to extract essential information from the CLIP (Radford et al., 2021; Dosovitskiy et al., 2020) image features. Alternatively, LLaVA (Liu et al., 2023c, b, 2024b) and MGM (Li et al., 2023d) utilize projection-based interfaces to facilitate interactions between text and image modalities. Furthermore, LLaMA-Adapter (Zhang et al., 2023a) and LaVIN (Luo et al., 2023) implement parameter-efficient tuning mechanisms to transfer image-related information to the LLM. A recent work also shows the usefulness of the object detection model for MLLMs (Jiao et al., 2024).
@@
From the perspective of datasets, there are two prevalent strategies: enhancing the quality of pretraining data and improving visual instruction tuning data. The former aims for better semantic alignment between images and text by introducing substantial numbers of image-text pairs, enabling MLLMs to proficiently address fundamental visual tasks, such as image captioning. However, MLLMs trained exclusively on pretraining datasets may encounter difficulties with VQA challenges. Recent research has increasingly concentrated on refining visual instruction tuning datasets, enabling MLLMs to enhance performance across diverse question-answering tasks. Works like LLaVA (Liu et al., 2023c, b, 2024b), InstructBLIP (Dai et al., 2023), SPHINX (Lin et al., 2023), and MGM (Li et al., 2023d) leverage high-quality finetuning datasets characterized by extensive task diversity, allowing models to excel in tasks related to image perception, reasoning, and optical character recognition (OCR). Additionally, methods such as Shikra (Chen et al., 2023), ASM (Wang et al., 2023), and PINK (Xuan et al., 2023) utilize substantial amounts of object detection data to enhance the models’ localization capabilities.
@@
In contrast to previous works, our research introduces a dataset that emphasizes image differences, showing empirical effectiveness and great potential to augment MLLMs’ VQA proficiency, object localization capabilities, and discernment of image distinctions.
Figure 2. An overview of the generation process for “object replacement” data.
[…]
An overview of the generation steps for “object replacement” data.
2.2. Datasets on Image Differences
@@
Datasets focused on image differences typically consist of pairs of similar images supplemented with textual descriptions of their variations. For instance, the Spot-the-Diff dataset (Jhamtani and Berg-Kirkpatrick, 2018) contains pairs of street scenes captured at different times by the same surveillance cameras. The CLEVR-Change dataset (Park et al., 2019) delineates scene variations of geometric objects against a clean backdrop. The Birds-to-Words dataset (Wah et al., 2011) elaborates on the nuanced differences among various bird species found in natural habitats. The Image-Edit-Request (Tan et al., 2019) dataset features edited images alongside their original counterparts, accompanied by descriptions of modifications made.
@@
Leveraging advancements in image editing technologies, some studies have employed generative models and editing techniques to create datasets centered on image differences. A prime example is InstructPix2Pix (Brooks et al., 2023), which utilizes the Prompt-to-Prompt (Hertz et al., 2022) image editing technique to direct Stable-Diffusion-1.5 (Rombach et al., 2022) in generating pairs of similar images, while employing GPT-3 (Mann et al., 2020) to craft the edited text as reference captions for image differences.
@@
Our approach, similar to InstructPix2Pix, employs the Prompt-to-Prompt technique alongside a generative model Stable-Diffusion-XL (Podell et al., 2023) to generate pairs of similar images. However, we incorporate multiple filtering stages to ensure data quality, with a particular emphasis on producing difference captions that focus on specific bounding box regions rather than the entire image, which ensures greater accuracy.
2.3. Models for Image Difference Captioning
@@
Image Difference Captioning (IDC) represents a specialized domain within image captioning characterized by its focus on subtle variations between images, rather than merely identifying common objects. The pioneering work in IDC, Spot-the-Diff (Jhamtani and Berg-Kirkpatrick, 2018), suggests potential change clusters and employs an LSTM (Hochreiter and Schmidhuber, 1997) decoder to generate captions addressing these variations. DUDA (Park et al., 2019) explores image differences at the semantic level using ResNet (He et al., 2016) and an LSTM to compute dynamic attention weights, producing captions that accurately depict these differences. VARD (Tu et al., 2023a) introduces a viewpoint-adaptive representation disentanglement network based on LSTM for differentiating between real and pseudo changes. Meanwhile, NCT (Tu et al., 2023b) employs a transformer (Vaswani et al., 2017) to integrate neighboring features, and CLIP4IDC (Guo et al., 2022) uses BERT-like training methodologies, adapting a CLIP model for IDC tasks.
With the emergence of MLLMs, VIXEN (Black et al., 2024) marks the inaugural use of these models for IDC tasks, attempting to extract features from CLIP outputs using linear layers or Q-formers and subsequently concatenating the features of differential images for input into an LLM to generate image difference captions.
@@
In contrast, our dataset is specifically designed for MLLMs. We construct our data based on the visual instruction tuning format established by mainstream MLLMs like LLaVA-1.5 and MGM, highlighting a new direction for exploration aimed at enhancing MLLMs from a data-centric perspective.
@@
Figure 3. An overview of the Difference Area Generator and its three internal components: Image Similarity Filter, Image-text Matching Filter, and Difference Detector.
[…]
An overview of the Difference Area Generator and its three internal components: Image Similarity Filter, Image-text Matching Filter, and Difference Detector.
3. The Curation of Img-Diff
3.1. Overview
@@
In recent years, contrastive learning have significantly enhanced the image-text understanding of vision-language models (Radford et al., 2021; Zhang et al., 2023b). These methods typically involve constructing batches of images and texts, requiring the model to distinguish between matching and non-matching image-text pairs, which improves its ability to differentiate between semantically similar and dissimilar pairs.
@@
Our approach leverages the principles of contrastive learning to generate MLLM image-text data. Specifically, it focuses on replacing objects within image pairs, encouraging MLLMs to identify similarities and differences in specific regions. This methodology aims to enhance the models’ capacity to recognize fine-grained differences in images, guided by textual descriptions that highlight detailed distinctions.
@@
The process of generating “object replacement” data can be divided into three key parts. The first part involves creating similar images and forming image pairs, where the only difference between these pairs is the object replacement (Section 3.2). The second part, referred to as the “Difference Area Generator”, extracts bounding box regions that contain object differences between the image pairs (Section 3.3). The third part, termed as the “Difference Captions Generator” (Section 3.4), utilizes MLLMs to generate descriptive text for the areas with object differences and creates question-answer pairs, such as “What objects have changed in this area?”
@@
This process incorporates multiple filtering operations, with specific thresholds outlined in Appendix B.4, which we determine through experimental comparisons in Section 4.6. Additional experimental details, including model selection and the training process, are also presented in Appendix B. For readability, we present several specific data examples in Appendix C and illustrate the general framework in Figure 2.
@@
To enhance reproducibility and reusability, our proposed components and end-to-end construction workflow are implemented as data processing operators and configurable data recipes within Data-Juicer (Chen et al., 2024a, b).
3.2. Image Pairs Generation
Firstly, we need to generate numerous pairs of similar images as candidates. The process is shown in Figure 2, which involves replacing the objects in images. We refer to InstructPix2Pix (Brooks et al., 2023) and employ a generative model combined with image editing techniques to generate image pairs. However, while InstructPix2Pix utilizes the Stable-Diffusion-1.5, we use the more advanced Stable-Diffusion-XL (Podell et al., 2023) to produce more realistic images.

### v3

@@
Abstract
High-performance Multimodal Large Language Models (MLLMs) are heavily dependent on data quality. To advance fine-grained image recognition within MLLMs, we introduce a novel data synthesis method inspired by contrastive learning and image difference captioning. Our key idea involves challenging the model to discern both matching and distinct elements by scrutinizing object differences in detailed regions across similar images. We begin by generating pairs of similar images that emphasize object variations. Following this, we employ a Difference Area Generator to pinpoint object differences, and subsequently, a Difference Captions Generator to articulate these differences. This process results in a high-quality dataset of ”object replacement” samples, termed Img-Diff, which can be scaled as needed due to its automated nature. We leverage this generated dataset to fine-tune state-of-the-art (SOTA) MLLMs, such as InternVL2, achieving substantial improvements across various image difference and Visual Question Answering tasks. Notably, the trained models significantly outperform existing SOTA models like GPT-4V and Gemini on the MMVP benchmark. Additionally, we conduct comprehensive evaluations to validate the dataset’s diversity, quality, and robustness, offering several insights into the synthesis of such contrastive datasets. We release our codes and dataset to encourage further research on multimodal data synthesis and MLLMs’ fundamental capabilities for image understanding.
@@ Abstract
Figure 1: Three “object replacement” examples within Img-Diff, highlighting fine-grined difference in both vision and language.
1 Introduction
@@
The emergence of large language models (LLMs) has revolutionized natural language processing [49, 64], also paved the way for the development of Multimodal Large Language Models (MLLMs) that seamlessly integrate linguistic and visual understanding. Improving the performance of MLLMs hinges on two primary avenues: evolving model architectures and enhancing dataset quality [54]. The majority of state-of-the-art (SOTA) MLLMs [43, 42, 44, 3, 38] employ a two-phase training strategy, commencing with a pre-training phase involving extensive image-text pairs for modality alignment, followed by a fine-tuning phase aimed at optimizing visual question answering (VQA) capabilities with specific instruction tuning datasets.
@@
The efficacy of pre-training datasets profoundly affects MLLMs’ capabilities in performing core visual tasks. Concurrently, the quality of visual instruction tuning datasets plays a crucial role in MLLMs’ overall performance in VQA tasks and diverse downstream applications. With the evolution of visual instruction tuning datasets, several recent studies have explored the integration of object detection and Optical Character Recognition (OCR) datasets, such as RefCOCO [29], Visual Genome [32], OCR-VQA [50], and TextVQA [60], effectively enhancing MLLMs’ proficiency in tasks requiring detailed image perception.
@@
In this paper, we focus on a new direction for enhancing MLLM datasets, driven by the potential of object variations in image pairs to refine models’ image recognition capabilities, as demonstrated by advancements in contrastive learning and image difference captioning [76, 5, 52, 26]. Specifically, we introduce a general-purpose yet challenging dataset named Img-Diff, which sets itself apart from existing visual instruction tuning datasets by generating pairs of highly similar images featuring subtle object alterations. Rather than compelling MLLMs to focus solely on a single image, our dataset challenges them to analyze paired images and articulate the differences within designated regions, meanwhile taking the high-quality textual descriptions of the difference as learning signals. By doing so, we aim to empower MLLMs with enhanced capabilities of fine-grained image recognition.
@@
To evaluate the effectiveness of our data synthesis method, we integrate the generated dataset into the original visual instruction tuning datasets of LLaVA-1.5 [42], MGM [38], and InternVL2 [12], and conduct fine-tuning. Subsequently, we evaluate their performance on image difference benchmarks, including MMVP [63], Spot-the-Diff [26], and Image-Edit-Request [62], as well as widely recognized MLLM benchmarks. Our evaluation results reveal that after fine-tuning with the Img-Diff dataset, the MLLMs achieve notable enhancements in image difference benchmarks, aligning their performance with state-of-the-art (SOTA) models. For instance, they notably surpass the SOTA models GPT-4V [1] and Gemini [16] on the MMVP benchmark, achieving an improvement of up to 12. Moreover, the models exhibit comprehensive improvements across eight well-recognized MLLM benchmarks, achieving an average score improvement of up to 3.06%, underscoring the useful role our dataset plays in bolstering MLLMs’ competencies in both image difference recognition and fine-grained image analysis.
@@
We further evaluate the diversity and quality of our dataset, ensuring it encompasses a broad array of object categories while showcasing rich variability. Through meticulous manual labeling, we affirm the high quality of our dataset. Furthermore, we conduct ablation studies to examine the effects of various filtering intensities. We also investigate an alternative methodology for constructing image difference data focusing on “object removal”, assessing its effectiveness and presenting fruitful insights on the construction of contrastive data.
@@
  • We present a novel data synthesis method and an effect-proven Img-Diff dataset, comprising pairs of highly similar images, with a focus on processes such as segmentation, filtering, and detailed captioning of image differences.
@@
  • We conduct visual instruction tuning on LLaVA-1.5-7B, MGM-7B, and InternVL2-8B using our dataset, and rigorously assess the fine-tuned models’ performance on many widely-used MLLM benchmarks and image difference benchmarks. Our dataset brings substantial performance improvements to the fine-tuned MLLMs.
@@
  • We provide a comprehensive evaluation of the diversity and quality of the generated dataset, confirming its richness and high standard. Through ablation studies, we identify good empirical filtering thresholds for such contrastive dataset.
@@
  • We open-source our dataset and codes at https://github.com/modelscope/data-juicer/tree/ImgDiff to facilitate ongoing research, encouraging innovative endeavors in MLLM datasets and image difference methods.
2 Background and Related Works
2.1 Multimodal Large Language Models
@@
Multimodal Large Language Models (MLLMs) have exhibited remarkable advancements since their introduction [1, 44, 45, 16, 38]. Research highlights two key factors that primarily influence the effectiveness of MLLMs: model architecture and dataset quality [54, 8].
@@
With respect to model architecture, notable approaches [2, 25, 33, 36, 3] leverage learnable queries to extract essential information from CLIP [55, 15] image features. Alternatively, LLaVA [43, 42, 44] and MGM [38] utilize projection-based interfaces to facilitate interactions between text and image modalities. Furthermore, LLaMA-Adapter [77] and LaVIN [48] implement parameter-efficient tuning mechanisms to transfer image-related information to the LLM. A recent work also show usefullness of object detection model for MLLMs [27].
@@
From the perspective of datasets, there are two prevalent strategies: enhancing the quality of pre-training data and improving visual instruction tuning data. The former aims for better semantic alignment between images and text by introducing abundant image-text pairs, enabling MLLMs to proficiently address fundamental visual tasks, such as image captioning. As for the latter, recent research has increasingly concentrated on refining visual instruction tuning datasets, enhancing MLLMs’ performance across various question-answering tasks. Works like LLaVA [43, 42, 44], SPHINX [41], MGM [38], and InternVL2 [12] leverage high-quality fine-tuning datasets characterized by extensive task diversity, allowing models to excel in tasks related to image perception, reasoning, and optical character recognition. Additionally, methods such as Shikra [10], ASM [70], and PINK [71] utilize substantial amounts of object detection data to enhance the models’ localization capabilities.
@@
In contrast to previous works, our research introduces a fully automated data synthesis method and generate a dataset that emphasizes image differences, showing empirical effectiveness and great potential to augment MLLMs’ VQA proficiency, object localization capabilities, and discernment of image distinctions.
2.2 Image Difference Datasets
@@
Datasets focused on image differences typically consist of pairs of similar images supplemented with textual descriptions of their variations. For instance, the Spot-the-Diff dataset [26] contains pairs of street scenes captured at different times by the same surveillance cameras. The CLEVR-Change dataset [52] delineates scene variations of geometric objects against a clean backdrop. The CUB-Bird dataset [69] focuses on the nuanced differences among various bird species found in natural habitats. The Image-Edit-Request [62] dataset features edited images, accompanied by descriptions of modifications made.
@@
Leveraging advancements in image editing technologies, some studies have employed generative models and editing techniques to create datasets centered on image differences. A prime example is InstructPix2Pix [6], which utilizes the Prompt-to-Prompt [20] image editing technique to direct Stable-Diffusion-1.5 [56] in generating pairs of similar images, while employing GPT-3 [49] to craft the edited text as reference captions for image differences.
@@
Our approach, referring to InstructPix2Pix, employs the Prompt-to-Prompt technique alongside an advanced generative model Stable-Diffusion-XL [53], which produces more realistic images, to generate pairs of similar images. Unlike InstructPix2Pix, we incorporate multiple filtering stages to ensure data quality, with a particular emphasis on producing difference captions that focus on specific regions rather than the entire image, which ensures greater accuracy.
2.3 Models for Image Difference Captioning
@@
Image Difference Captioning (IDC) represents a specialized domain within image captioning characterized by its focus on subtle variations between images. As for the pioneering work in IDC, Spot-the-Diff [26] presents potential change clusters and employs an LSTM [21] decoder to generate difference captions. DUDA [52] explores image differences at the semantic level, using a ResNet [19] and an LSTM to compute dynamic attention weights and produce captions. VARD [65] introduces a viewpoint-adaptive representation disentanglement network based on LSTM for differentiating between real and pseudo changes. Meanwhile, NCT [66] employs a transformer [67] to integrate neighboring features, and CLIP4IDC [18] uses BERT-like training methodologies, adapting a CLIP model for IDC tasks. With the emergence of MLLMs, VIXEN [5] marks the inaugural use of these models for IDC tasks, mapping the features of differential images to text space and employing an LLM to generate image difference captions.
@@
Our data synthesis method and dataset is specifically designed for MLLMs. We build our data in accordance with the instruction-following format established by mainstream MLLMs such as LLaVA-1.5 and MGM, highlighting a new direction for exploration aimed at enhancing MLLMs from a data-centric perspective.
Figure 2: The generation process for “object replacement” data.
@@
Figure 3: An overview of the Difference Area Generator and its three internal components: Image Similarity Filter, Image-text Matching Filter, and Difference Detector.
3 Methodology
3.1 Overview
@@
In recent years, contrastive learning have significantly enhanced the image-text understanding of vision-language models [55, 76]. These methods typically involve constructing batches of images and texts, requiring the model to distinguish between matching and non-matching image-text pairs, which improves its ability to differentiate between semantically similar and dissimilar pairs.
@@
Our method leverages the principles of contrastive learning to generate MLLM image-text data. Specifically, it focuses on replacing objects within image pairs, encouraging MLLMs to identify similarities and differences in specific regions. This method aims to enhance models’ capacity to recognize fine-grained differences in images, guided by textual descriptions that highlight detailed distinctions.
@@
As shown in Figure 2, the process of generating “object replacement” data can be divided into three key parts. The first part involves creating similar images and forming image pairs, where the only difference between the images in pairs is the objects replaced (Section 3.2). The second part, named the “Difference Area Generator”, extracts bounding box regions that contain object differences between the images in pairs (Section 3.3). The third part, termed as the “Difference Captions Generator” (Section 3.4), utilizes an MLLM to generate descriptive text for the areas with object differences and creates question-answer pairs.
@@
This process incorporates multiple filtering operations, with specific thresholds outlined in Section 15.4, which we determine through experimental comparisons in Section 14. Additional experimental details, including model selection and the time consumption, are also presented in Section 15. For readability, we present data examples in Section 17.
@@
To enhance reproducibility and reusability, our proposed components and end-to-end construction workflow are implemented as data processing operators and configurable data recipes within Data-Juicer [7, 9].
3.2 Image Pairs Generation
The first step of our data synthesis method is to generate pairs of similar images as candidates. The process is shown in Figure 2. We employ a generative model called Stable-Diffusion-XL [53] and an image editing technique called Prompt-to-Prompt [20] to generate image pairs that highlight object replacement.